# A Temporal-Spectral Fusion Transformer with Subject-specific Adapter for Enhancing RSVP-BCI Decoding

### v1

@@ Abstract
Brain-computer Interface (BCI), Rapid Serial Visual Presentation (RSVP), Multi-view Learning, Adapter-based Fine-tuning, Transformer.
@@ 1 Introduction
The Brain-computer Interface (BCI) systems facilitate the translation of neural signals into computer commands, establishing a direct pathway for communication between the human brain and external devices wolpaw2002brain . In BCI systems, electroencephalography (EEG) is generally utilized owing to its high temporal resolution, cost-effectiveness, portability, and noninvasive applicability nicolas2012brain ; ang2013brain . Rapid Serial Visual Presentation (RSVP)-based BCI systems have attracted significant attention due to their potential for enhancing human-computer collaboration cecotti2015single ; marathe2015effect ; lees2018review and have been applied in various fields, including target image retrieval cecotti2010convolutional ; alpert2013spatiotemporal ; marathe2015improved , spellers acqualagna2013gaze , anomaly detection barngrover2015brain , and anti-deception wu2018anti .
In the RSVP-BCI system, an image sequence including various classes is presented to a subject at a high rate cecotti2014single . The subject needs to identify the specific target images from other distractors (i.e. nontarget images), where the target images are sparsely distributed in a random order. As a result, these rare target images evoke event-related potentials (ERPs) containing the P300 component which is a prominent positive peak typically observed around 300 ms after infrequent stimuli squires1976effect ; polich2007updating . The retrieval of the target images can be realized through the detection of the P300 component within EEG signals (see Fig. 1).
@@
P300 components in the same task are relatively consistent in terms of both amplitude and latency, but it is still challenging to achieve single-trial RSVP decoding. Many algorithms have been developed based on conventional machine learning methods, which are applicable to RSVP decoding. In 2006, Gerson et al. proposed a linear discriminant method called Hierarchical Discriminant Component Analysis (HDCA) which operates on channel and time dimensions of EEG signals HDCA . Barachant et al. introduced Riemann geometry to measure the distance between EEG raw signals and proposed the Minimum Distance to Riemann Mean (MDRM) MDRM . These methods achieve RSVP decoding by leveraging the temporal domain characteristics of EEG temporal signals. Since ERPs can induce phase-synchronized oscillations in EEG signals, resulting in event-related changes in both amplitude and phase at specific frequencies makeig2002dynamic ; makeig2004mining , researchers also utilize the time-frequency characteristics of EEG signals for P300 detection robbins2020sensitive . Several methods have employed Continuous Wavelet Transform (CWT) to convert EEG temporal signals into spectrogram images for extracting spectral features and subsequently classified them using either Linear Discriminant Analysis (LDA) or Support Vector Machine (SVM) bostanov2004bci ; roach2008event .
@@
With the remarkable progress achieved by deep learning in numerous visual tasks krizhevsky2012imagenet ; lecun2015deep , deep learning has also been demonstrated to be useful in decoding EEG signals of RSVP tasks. Many deep learning methods, including MCNN MCNN , EEGNet EEGNet , One Convolution Layer Neural Network (OCLNN) OCLNN , Phase-locked Network (PLNet) PLNet , and Phase Preservation Neural Network (PPNN) PPNN , have improved the RSVP decoding performance by distinguishing temporal domain characteristics of EEG signals. Moreover, several methods that utilize EEG time-frequency information have been developed. For instance, Tajmirriahi et al. proposed an interpretable Convolutional Neural Network (CNN) for P300 detection by leveraging time-frequency characteristics tajmirriahi2022interpretable . Havaei et al. also incorporated the time-frequency localization properties of the Gabor transform and CNN to distinguish P300 signals havaei2023efficient . Because EEG signals can vary significantly across subjects and over time arising from factors such as mental states, the spatial geometry of the head, and the diversity of brain functions wolpaw2002brain ; morioka2015learning , the aforementioned methods require the collection of labeled data from each new subject on the same day to train individual models, which can be called the subject-dependent decoding.
@@
Enhancing the performance of deep learning models necessitates an extensive training dataset li2022tff ; wei2022erp . However, collecting a large amount of training data from the new subject before using the BCI system requires a time-consuming preparation procedure, which can delay the application of the RSVP-BCI system and lead to subject fatigue. This limits the potential for performance improvement in RSVP-BCI systems. To further improve RSVP decoding performance, researchers have attempted to use labeled data collected from existing subjects and a small amount of data from the new subject to jointly train the model to obtain better RSVP decoding performance on the new subject wei2020reducing ; fan2022dc . Wei et al. introduced a conditional adversarial network that aligns the feature distribution of each existing subject with that of the new subject, enabling the feature extractor trained on data from existing subjects to extract task-related features from data of the new subject wei2020reducing . Similarly, Fan et al. utilized the adversarial network to mitigate the distribution discrepancy between existing subjects and the new subject fan2022dc . These methods utilize the limited training data from new subjects to achieve better performance than comparison methods, thus reducing the requirement for training data from new subjects.
@@
In this paper, to enhance RSVP decoding performance while reducing preparation time for RSVP-BCI systems, we propose a Temporal-Spectral fusion transformer with Subject-specific Adapter (TSformer-SA) which is optimized through a two-stage training strategy. Inspired by studies that incorporate multi-view information in EEG decoding and obtain more discriminative high-level representations tian2019deep ; sun2022dual , we integrate the temporal and spectral information of EEG signals within a multi-view learning network based on Transformer, which is compatible with diverse modalities and can realize cross-view interaction vaswani2017attention ; tsai2019multimodal ; xu2023multimodal . The EEG temporal signals representing the temporal view and spectrogram images obtained through CWT representing the spectral view are utilized for input. Firstly, we propose a feature extractor to tokenize the input and extract global features within each view. Secondly, the cross-view interaction module and the multi-view consistency loss are designed to reduce the disparity between the two-view features from the same EEG signal for extracting common task-related representations. Thirdly, we propose an attention-based fusion module to effectively fuse the features extracted from two views to obtain discriminative features containing both temporal and spectral characteristics for classification. Moreover, the adapter-based fine-tuning can transfer learned general patterns of pre-trained models to downstream tasks and consequently enhance decoding performance by updating a few parameters houlsby2019parameter . Thus, instead of training the model from scratch with substantial quantities of existing data, we propose the subject-specific adapter and design a two-stage training approach to leverage the data from existing subjects while reducing the preparation time. In the two-stage training approach, we employ the data from existing subjects to pre-train the TSformer in advance and only fine-tune the subject-specific adapter using data from the new subject in the preparation procedure (see Fig. 2). The subject-specific adapter is utilized to adapt the model pre-trained on the data from existing subjects to the data characteristics of new subjects and enhance decoding performance with limited training data from new subjects. The main contributions of this work are summarized as follows:
@@
    We propose a novel Temporal-Spectral fusion transformer with Subject-specific Adapter (TSformer-SA) optimized through a two-stage training strategy to enhance RSVP decoding performance and reduce the preparation time of BCI systems. Codes will be released as soon as this paper is accepted for publication.
@@
    Extensive experiments are conducted on our self-collection open-source dataset (the link will be updated) to verify the performance of TSformer-SA. The experimental results demonstrate that our model achieves excellent decoding performance while reducing the requirement for new subjects’ training data and the training time in the preparation procedure.
@@ 2.1.1 Conventional Machine Learning Methods
In 2006, Gerson, A.D. et al. introduced HDCA, which employed linear discriminant analysis to compute spatial weighting vectors for various time intervals and project single-trial EEG data. This method has been widely adopted in RSVP decoding for its simplicity and efficiency HDCA . In 2014, Barachan et al. proposed the MDRM method MDRM , which was rooted in Riemannian geometry and classified covariance matrices derived from original EEG signals based on their minimum distance to the mean. In 2016, Waytowich et al. introduced Spectral Transfer Learning using Information Geometry (STIG) for subject-independent RSVP decoding, which was also based on Riemannian geometry. The STIG employed spectral-meta learning to ensemble information geometry classifiers trained on data from other training subjects for classification waytowich2016spectral .
@@ 2.1.2 CNN-based Methods
In recent years, deep learning methods have gained increasing popularity for classifying EEG data and demonstrated superior decoding performance. In 2011, Cecotti et al. first introduced the CNN for P300 detection. The network incorporated spatial and temporal convolution layers for feature extraction and included multiple model variants, which proves the effectiveness of CNN for RSVP decoding cecotti2010convolutional . Subsequently, Manor et al. (2015) applied CNN to the RSVP target image retrieval task and proposed a larger network combining a novel spatiotemporal regularization method specifically designed for EEG signals MCNN . In 2018, Lawhern Vernon J et al. introduced EEGNet, a compact CNN incorporating depthwise separable convolution chollet2017xception which can simplify the structure and reduce the number of parameters EEGNet . In 2020, Lee et al. introduced a model based on EEGNet for the subject-independent P300 BCI speller, which demonstrated comparable performance to subject-dependent methods LeeCNN . In 2021, Zang et al. leveraged the phase-locked characteristic of ERPs and proposed PLNet which separately processes spatial convolution in different time periods to extract spatiotemporal features for RSVP decoding. PLNet . Meanwhile, Li et al. proposed a phase preservation neural network (PPNN) to learn phase information of EEG signals in all frequency bands and improve the classification performance in the RSVP task PPNN .
@@ 2.2 Transformer-based EEG Decoding Methods
In comparison to convolutions which utilize identical filter weights for all input patches, Transformer is an attention-based structure that can flexibly adjust the receptive field and effectively capture global interactions via self-attention mechanism ramachandran2019stand ; naseer2021intriguing . Over the past two years, researchers have explored the use of Transformers in EEG decoding. In the emotion recognition task, Wang et al. proposed the hierarchical spatial learning transformer (HSLT) to extract discriminative features by capturing EEG spatial dependencies from electrode level to brain-region level wang . In 2022, Casal et al. employed a temporal convolutional network to extract original features as the input of the Transformer for learning the transition rules among sleep stages TCN-T . This model can be applied to both heart rate and EEG decoding tasks li2022tff . For SSVEP classification, Chen et al. (2023) proposed the SSVEPformer chen2023transformer , which leveraged complex spectrum features extracted from SSVEP data as input and explored spectral and spatial information. In Motor Imagery (MI) decoding, Song et al. (2023) proposed the EEG-conformer, a compact convolutional Transformer capable of unifying local and global features within a single EEG classification framework song2022eeg . All of these studies demonstrate the effectiveness of Transformers in EEG decoding.
@@ 2.3 Multi-view Learning EEG Decoding Methods
Multi-view representation learning tackles the challenge of learning effective representations from multi-view data for enhancing the development of prediction models li2018survey . In 2021, Li et al. proposed a novel temporal-spectral-based squeeze-and-excitation feature fusion network, which effectively fused temporal features extracted by temporal convolutions and spectral features extracted by multi-level wavelet convolutions for decoding MI-EEG signals li2021temporal . In 2022, Sun et al. proposed a model based on dynamic graph convolution and adaptive feature fusion sun2022dual . This model computed both differential entropy and power spectral density features from each EEG segment as two distinct views, and a dual-branch graph dynamic convolution (DBGC) network was designed to capture deeper temporal and spectral features. The next year, Luo et al. proposed a multi-view fusion model based on Transformer for EEG-based visual recognition luo2023dual , which fused features from the spatio-temporal and spectral-temporal domains of EEG signals and achieved superior performance compared to state-of-the-art methods. These approaches demonstrate that utilizing multi-view information of EEG signals enables the model to extract more discriminative high-level features, which enhances the classification performance.
@@ 3.1 Subjects
In the experiment, we design and implement three RSVP tasks for target image retrieval. The targets of the three tasks are planes, cars, and people, respectively. The Task plane involves 20 subjects (15 males and 5 females, aged 23.4 ± 1.1, with 18 being right-handed). The Task car involves 20 subjects (11 males and 9 females, aged 23.75 ± 1.3, all of whom are right-handed). The Task people involves 31 subjects (19 males and 12 females, aged 24.9 ± 2.8, with 28 being right-handed). There is no overlap in the subject participation across the three tasks. All participants possess normal or corrected-to-normal vision. Before the experiment, the Institutional Review Board of the Institute of Automation, Chinese Academy of Sciences, provides ethical approval. Additionally, all subjects are voluntary and provide written informed consent.
3.2 RSVP Paradigm
The target image retrieval experiment consists of three distinct tasks: Task plane, Task car, and Task people. The Task plane requires participants to identify the image that contains the target plane from remote sensing images. Participants in the Task car need to identify the image that includes the target car from drone-captured ground images. Task people requires participants to identify the images containing the target people from street scene images. The visual stimulus images used in Task plane, Task car, and Task people are from three different sources: the Dior dataset li2020object , our self-collection drone aerial images, and the scenes and objects database torralba2009csail , respectively. Within each task, the visual stimulus images that contain the task-specified target are designated as target images, while the remaining images are considered as nontarget images. Figure 3 (a) illustrates a few examples of the stimulus images.
Figure 3: Illustration of the RSVP paradigm. (a) Examples of target and nontarget images in Task plane, Task car, and Task people. (b) Experimental settings about the division of blocks and sequences for each subject.
The images are randomly presented at a frequency of 10 Hz, with the target images appearing with a probability of 4%. Participants are seated 1 meter away from the monitor with a resolution of 1920 x 1080. When participants achieve a state of calmness, they are presented with a sequence of images and instructed to identify the target images. As shown in Fig. 3 (b), each subject completes a task consisting of $m_{blk}$ blocks (Task plane: $m_{blk}=10$; Task car: $m_{blk}=5$; Task people: $m_{blk}=10$). Each block contains $m_{seq}$ sequences (Task plane: $m_{seq}=14$; Task car: $m_{seq}=16$; Task people: $m_{seq}=14$), where each sequence consists of 100 images. The interval between sequences is controlled by the participants. Each block takes around 5 minutes to complete, with a break time of approximately 2-3 minutes between blocks.

### v2

@@ Abstract
Brain-Computer Interface (BCI) , Rapid Serial Visual Presentation (RSVP) , Multi-View Learning , Transformer , Adapter-Based Fine-Tuning.
journal: Neural Networksfn0fn0footnotetext: The first two authors contributed equally to this work.
@@ 1 Introduction
The Brain-Computer Interface (BCI) systems facilitate the translation of neural signals into computer commands, establishing a direct pathway for communication between the human brain and external devices [1]. In BCI systems, electroencephalography (EEG) is generally utilized owing to its high temporal resolution, cost-effectiveness, portability, and noninvasive applicability [2, 3]. Rapid Serial Visual Presentation (RSVP)-based BCI systems have attracted significant attention due to their potential for enhancing human-computer collaboration [4, 5] and have been applied in various fields, including target image retrieval [6, 7], spellers [8], face recognition [9], anomaly detection [10], and anti-deception [11].
In the RSVP-BCI system, an image sequence including various classes is presented to a subject at a high rate [12]. The subject needs to identify the specific target images from other distractors (i.e. nontarget images), where the target images are sparsely distributed in a random order. As a result, these rare target images evoke event-related potentials (ERPs) containing the P300 component which is a prominent positive peak typically observed around 300 ms after infrequent stimuli [13, 14]. The retrieval of the target images can be realized through the detection of the P300 component within EEG signals (see Fig. 1).
@@
P300 components in the same task are relatively consistent in terms of both amplitude and latency, but it is still challenging to achieve single-trial RSVP decoding. Many algorithms have been developed based on conventional machine learning methods, which are applicable to RSVP decoding. In 2006, Gerson et al. proposed a linear discriminant method called Hierarchical Discriminant Component Analysis (HDCA) which operates on channel and time dimensions of EEG signals [15]. Barachant et al. introduced Riemann geometry to measure the distance between EEG raw signals and proposed the Minimum Distance to Riemann Mean (MDRM) [16]. These methods achieve RSVP decoding by leveraging the temporal domain characteristics of EEG temporal signals. Since ERPs can induce phase-synchronized oscillations in EEG signals, resulting in event-related changes in both amplitude and phase at specific frequencies [17, 18], researchers also utilize the time-frequency characteristics of EEG signals for P300 detection [19]. Several methods have employed Continuous Wavelet Transform (CWT) to convert EEG temporal signals into spectrogram images for extracting spectral features and subsequently classified them using either Linear Discriminant Analysis (LDA) or Support Vector Machine (SVM) [20, 21].
@@
With the remarkable progress achieved by deep learning in numerous visual tasks [22, 23], deep learning has also been demonstrated to be useful in decoding EEG signals of RSVP tasks. Many deep learning methods, including MCNN [24], EEGNet [25], One Convolution Layer Neural Network (OCLNN) [26], Phase-locked Network (PLNet) [27], and Phase Preservation Neural Network (PPNN) [28], have improved the RSVP decoding performance by distinguishing temporal domain characteristics of EEG signals. Moreover, several methods that utilize EEG time-frequency information have been developed. For instance, Tajmirriahi et al. proposed an interpretable Convolutional Neural Network (CNN) for P300 detection by leveraging time-frequency characteristics [29]. Havaei et al. also incorporated the time-frequency localization properties of the Gabor transform and CNN to distinguish P300 signals [30]. Because EEG signals can vary significantly across subjects and over time arising from factors such as mental states, the spatial geometry of the head, and the diversity of brain functions [31], the aforementioned methods require the collection of labeled data from each new subject on the same day to train individual models, which can be called the subject-dependent decoding.
@@
Enhancing the performance of deep learning models necessitates an extensive training dataset [32]. However, collecting a large amount of training data from the new subject before using the BCI system requires a time-consuming preparation procedure, which can delay the application of the RSVP-BCI system and lead to subject fatigue. This limits the potential for performance improvement in RSVP-BCI systems. To further improve RSVP decoding performance, researchers have attempted to use labeled data collected from existing subjects and a small amount of data from the new subject to jointly train the model to obtain better RSVP decoding performance on the new subject [33]. Wei et al. introduced a conditional adversarial network that aligns the feature distribution of each existing subject with that of the new subject, enabling the feature extractor trained on data from existing subjects to extract task-related features from data of the new subject [33]. Similarly, Fan et al. utilized the adversarial network to mitigate the distribution discrepancy between existing subjects and the new subject [34]. These methods utilize the limited training data from new subjects to achieve better performance than comparison methods, thus reducing the requirement for training data from new subjects.
@@
In this paper, to enhance RSVP decoding performance while reducing preparation time for RSVP-BCI systems, we propose a Temporal-Spectral fusion transformer with Subject-specific Adapter (TSformer-SA) which is optimized through a two-stage training strategy. Inspired by studies that incorporate multi-view information in EEG decoding and obtain more discriminative high-level representations [35, 36, 37], we integrate the temporal and spectral information of EEG signals within a multi-view learning network based on Transformer, which is compatible with diverse modalities and can realize cross-view interaction [38]. The EEG temporal signals representing the temporal view and spectrogram images obtained through CWT representing the spectral view are utilized for input. Firstly, we propose a feature extractor to tokenize the input and extract global features within each view. Secondly, the cross-view interaction module and the multi-view consistency loss are designed to reduce the disparity between the two-view features from the same EEG signal for extracting common task-related representations. Thirdly, we propose an attention-based fusion module to effectively fuse the features extracted from two views to obtain discriminative features containing both temporal and spectral characteristics for classification. Moreover, the adapter-based fine-tuning can transfer learned general patterns of pre-trained models to downstream tasks and consequently enhance decoding performance by updating a few parameters [39]. Thus, instead of training the model from scratch with substantial quantities of existing data, we propose the subject-specific adapter and design a two-stage training approach to leverage the data from existing subjects while reducing the preparation time. In the two-stage training approach, we employ the data from existing subjects to pre-train the TSformer in advance and only fine-tune the subject-specific adapter using data from the new subject in the preparation procedure (see Fig. 2). The subject-specific adapter is utilized to adapt the model pre-trained on the data from existing subjects to the data characteristics of new subjects and enhance decoding performance with limited training data from new subjects. The main contributions of this work are summarized as follows:
@@
    We propose a novel Temporal-Spectral fusion transformer with Subject-specific Adapter (TSformer-SA) optimized through a two-stage training strategy to enhance RSVP decoding performance and reduce the preparation time of BCI systems. The codes have been released at: https://github.com/lixujin99/TSformer-SA.
@@
    Extensive experiments are conducted on our self-collection open-source dataset to verify the performance of TSformer-SA. The experimental results demonstrate that our model achieves excellent decoding performance while reducing the requirement for new subjects’ training data and the training time in the preparation procedure.
@@ 2.1.1 Conventional Machine Learning Methods
In 2006, Gerson, A.D. et al. introduced HDCA, which employed linear discriminant analysis to compute spatial weighting vectors for various time intervals and project single-trial EEG data. This method has been widely adopted in RSVP decoding for its simplicity and efficiency [15]. In 2014, Barachan et al. proposed the MDRM method [16], which was rooted in Riemannian geometry and classified covariance matrices derived from original EEG signals based on their minimum distance to the mean. In 2016, Waytowich et al. introduced Spectral Transfer Learning using Information Geometry (STIG) for subject-independent RSVP decoding, which was also based on Riemannian geometry. The STIG employed spectral-meta learning to ensemble information geometry classifiers trained on data from other training subjects for classification [40].
@@ 2.1.2 CNN-Based Methods
In recent years, deep learning methods have gained increasing popularity for classifying EEG data and demonstrated superior decoding performance. In 2011, Cecotti et al. first introduced the CNN for P300 detection. The network incorporated spatial and temporal convolution layers for feature extraction and included multiple model variants, which proves the effectiveness of CNN for RSVP decoding [41]. Subsequently, Manor et al. (2015) applied CNN to the RSVP target image retrieval task and proposed a larger network combining a novel spatiotemporal regularization method specifically designed for EEG signals [24]. In 2018, Lawhern Vernon J et al. introduced EEGNet, a compact CNN incorporating depthwise separable convolution [42] which can simplify the structure and reduce the number of parameters [25]. In 2020, Lee et al. introduced a model based on EEGNet for the subject-independent P300 BCI speller, which demonstrated comparable performance to subject-dependent methods [43]. In 2021, Zang et al. leveraged the phase-locked characteristic of ERPs and proposed PLNet which separately processes spatial convolution in different time periods to extract spatiotemporal features for RSVP decoding. [27]. Meanwhile, Li et al. proposed a phase preservation neural network (PPNN) to learn phase information of EEG signals in all frequency bands and improve the classification performance in the RSVP task [28].
@@ 2.2 Transformer-Based EEG Decoding Methods
In comparison to convolutions which utilize identical filter weights for all input patches, Transformer is an attention-based structure that can flexibly adjust the receptive field and effectively capture global interactions via self-attention mechanism [44, 45]. Over the past two years, researchers have explored the use of Transformers in EEG decoding. For SSVEP classification, Chen et al. (2023) proposed the SSVEPformer [46], which leveraged complex spectrum features extracted from SSVEP data as input and explored spectral and spatial information. In Motor Imagery (MI) decoding, Song et al. (2023) proposed the EEG-conformer, a compact convolutional Transformer capable of unifying local and global features within a single EEG classification framework [47]. In the emotion recognition task, Casal et al. (2022) employed a temporal convolution network to extract original features as the input of the Transformer for learning the transition rules among sleep stages [48]. This model can be applied to both heart rate and EEG decoding tasks [32]. In 2024, Yan et al. combined a graph convolution network and multi-scale transformer for EEG emotion recognition, where the graph-based networks primarily focused on local EEG channel relationships and the multi-scale transformer learned spatial features more elaborately to enhance feature representation ability [49]. All of these studies demonstrate the effectiveness of Transformers in EEG decoding.
@@ 2.3 Multi-View Learning EEG Decoding Methods
Multi-view representation learning tackles the challenge of learning effective representations from multi-view data for enhancing the development of prediction models [50]. In 2021, Li et al. proposed a novel temporal-spectral-based squeeze-and-excitation feature fusion network, which effectively fused temporal features extracted by temporal convolutions and spectral features extracted by multi-level wavelet convolutions for decoding MI-EEG signals [51]. The next year, Sun et al. proposed a model based on dynamic graph convolution and adaptive feature fusion [36]. This model computed both differential entropy and power spectral density features from each EEG segment as two distinct views, and a dual-branch graph dynamic convolution (DBGC) network was designed to capture deeper temporal and spectral features. In 2023, Luo et al. proposed a multi-view fusion model based on Transformer for EEG-based visual recognition [52], which fused features from the spatio-temporal and spectral-temporal domains of EEG signals and achieved superior performance compared to state-of-the-art methods. These approaches demonstrate that utilizing multi-view information of EEG signals enables the model to extract more discriminative high-level features, which enhances the classification performance.
@@ 3 Materials
3.1 RSVP Paradigm
In this study, we aim to propose a method to enhance the RSVP decoding performance and evaluate its performance in different practical application scenarios. Therefore, we design three different RSVP experiments to conduct three independent datasets.
3.1.1 RSVP Tasks
Figure 3: Illustration of the RSVP paradigm. (a) Examples of target and nontarget images in Task plane, Task car, and Task people. The stimulus images in Task plane are sourced from the remote sensing Dior dataset [53], the stimulus images of the Task car are from our self-collection drone aerial images, and the stimulus images of the Task people are from the scenes and objects database [54]. (b) Experimental settings about the division of blocks and sequences for each subject.
At present, retrieving images containing target people from street view images is the most commonly used RSVP task in previous studies [41, 6, 7, 55, 56], which has been verified the validity of this paradigm [40]. Hence, in this study, we select the scenes and objects database [54] to conduct an RSVP task that retrieves target people from street view images to evaluate the model performance. To enhance the diversity of RSVP tasks in our study and focus on practical applications, we also design two RSVP tasks to simulate optical remote sensing observation [57, 58] and drone observation [59] respectively. We utilize the remote sensing Dior dataset [53] and our self-collection drone aerial images to simulate observation, respectively. Considering that the majority of our subjects are college students, we choose common objects that are familiar to the subjects and have high practical value as targets, so that the subjects can start the target retrieval task with simple training duration. Hence, in the remote sensing observation, we select the images containing a plane in the Dior dataset as the target, and we select the car as the target of drone observation in the drone aerial images.
As a result, we have designed three distinct RSVP tasks: (1) Task plane involves identifying planes from remote sensing images, which is designed to simulate remote sensing observation; (2) Task car focuses on identifying cars from images captured by drones, which simulates the drone observation; (3) Task people involves retrieving images containing people from street scene images, which simulates the security monitor in street view. Within each task, the visual stimulus images that contain the task-specified target are designated as target images, while the remaining images are considered nontarget images. Figure 3 (a) illustrates a few examples of the stimulus images. Finally, we recruit participants for each of the three RSVP tasks respectively.
3.1.2 Parameters of RSVP Paradigm
The images in each RSVP task are randomly presented at a frequency of 10 Hz, with the target images appearing with a probability of 4%. Participants are seated 1 meter away from the monitor with a resolution of 1920 x 1080. When participants achieve a state of calmness, they are presented with a sequence of images and instructed to identify the target images. As shown in Fig. 3 (b), each subject completes a task consisting of $m_{blk}$ blocks (Task plane: $m_{blk}=10$; Task car: $m_{blk}=5$; Task people: $m_{blk}=10$). Each block contains $m_{seq}$ sequences (Task plane: $m_{seq}=14$; Task car: $m_{seq}=16$; Task people: $m_{seq}=14$), where each sequence consists of 100 images. The interval between sequences is controlled by the participants. Each block takes around 5 minutes to complete, with a break time of approximately 2-3 minutes between blocks.
@@ 3.2 Subjects
In the experiment, we design and implement three RSVP tasks for target image retrieval including Task plane, Task car, and Task people. Each task is an independent dataset. The Task plane involves 20 subjects (15 males and 5 females, aged 23.4 ± 1.1, with 18 being right-handed). The Task car involves 20 subjects (11 males and 9 females, aged 23.75 ± 1.3, all of whom are right-handed). The Task people involves 31 subjects (19 males and 12 females, aged 24.9 ± 2.8, with 28 being right-handed).
There is no overlap in the subject participation across the three tasks. All participants possess normal or corrected-to-normal vision. Before the experiment, the Institutional Review Board of the Institute of Automation, Chinese Academy of Sciences, provides ethical approval. Additionally, all subjects are voluntary and provide written informed consent.